When air (21 mol% O2, 79 mol% N2) is placed in contact with 1000 cm3 of liquid water at body temperature, 37°C, and 1 atm absolute, approximately 14.1 standard cubic centimeters [cm3(STP)] of gas are absorbed in the water at equilibrium. Subsequent analysis of the liquid reveals that 33.4 mol% of the dissolved gas is oxygen and the balance is oxygen

a) Estimate the Henry’s law coefficients (atm/mole fraction) of oxygen and nitrogen at 37°C

b) An adult absorbs approximately 0.4g O2/min in the blood flowing through the lungs. Assuming that blood behaves like water and that it enters the lungs free of oxygen, estimate the flow rate of blood into the lungs in L/min. (answer: 59 L blood/min)
Answer #1

a) Henry's law is P = KH X

KH = P/X

For oxygen KH = P/X = 1/ 21 = 0.0476 atm/mol

For nitrogen KH = P/X = 1/79 = 0.01266 atm/mol